OpenStudy (mathmath333):

If xy is a positive 2-digit number and u, v, x, y are digits, then find the number of solutions of the question & (xy)^{2}=u!+v

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{If xy is a positive 2-digit number and u, v, x, y are digits,}\hspace{.33em}\\~\\ & \normalsize \text{ then find the number of solutions of the question}\hspace{.33em}\\~\\ & (xy)^{2}=u!+v \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

i would say at most 7, since 8! >>> 99^2 ^^

OpenStudy (anonymous):

hey, I found one 71^2 = 7! + 1 =]]

OpenStudy (anonymous):

oh hey I found yet another one 27^2 = 6! + 9 =]]

OpenStudy (anonymous):

12^2 = 5! + 24 =]]]

OpenStudy (anonymous):

7^2 = (4! + 25), but 7 is a one digit number. So not part of the solution Answer: 3 solutions ^^

OpenStudy (mathmath333):

how did u find that

OpenStudy (anonymous):

trial and error XD. I started with 7! and added numbers to see if I have a perfect square.